OpenStudy (anonymous):

Is the real number sqrt1.21 rational or irrational?

OpenStudy (anonymous):



OpenStudy (mosaic):

First find the square root of 1.21. What is it?

OpenStudy (john_es):

You can rewrite 1.21 with its rational form, \[1.21=\frac{121}{100}\]Now you can ask if \[\sqrt{\frac{121}{100}}\] is rational or irrational.

OpenStudy (anonymous):

ok, sqrt of 1.21 is 1.1 i think

OpenStudy (anonymous):

making it rational.

OpenStudy (mosaic):

Yes, 1.1. And 1.1 can be written as 11/10. The definition of a rational number is that it can be written as a ratio of two integers.

OpenStudy (mosaic):

Yes, it is rational.